OpenStudy (anonymous):

the perimeter of a triangle is 76 cm. side a of the triangle is twice as long as side b. side c is 1 cm longer than side a. find the length of each side.

OpenStudy (anonymous):

hmm do u want the answer or just a hint?

OpenStudy (anonymous):

the answer

OpenStudy (anonymous):

 x+2x+2x+1=76 5x = 76-1 x=15